### v2

# Characterising Ball Quotients through their (higher) Chern Numbers

Abstract.
In this short note we provide a characterisation of ball quotients among all minimal smooth projective varieties of general type purely in terms of their characteristic numbers. This generalises earlier work of Miyaoka, Yau and Greb–Kebekus–Peternell–Taji.
Key words and phrases:
Uniformisation, Chern classes, minimal varieties, Hermitian symmetric domains
2020 Mathematics Subject Classification:
Primary 32Q30, Secondary 14E30, 32Q45.
1. Introduction
A ball quotient variety is a compact complex manifold $X$ whose universal cover is biholomorphic to the complex unit ball $\mathbb{B}^{n}=\{z\in\mathbb{C}^{n}\ |\ \|z\|^{2}<1\}$. In other words, $X\cong\mathbb{B}^{n}/\Lambda$, where $\Lambda\subseteq\mathrm{Aut}(\mathbb{B}^{n})=\mathrm{PU}(n,1)$ is a torsion-free and cocompact lattice. In this case, it is well-known that the canonical line bundle $\omega_{X}$ is positive; in particular, ball quotient varieties are always projective.
A famous source of examples of ball quotients are fake projective planes, i.e. smooth projective surfaces $S$ which are not isomorphic to $\mathbb{P}^{2}$ but satisfy $H^{i}(S,\mathbb{Q})\cong H^{i}(\mathbb{P}^{2},\mathbb{Q})$; see [PrasadYeung_FakeProjectivePlanes, CartwrightSteger_EnumerationOfAllFakeProjectivePlanes] for a complete classification of the corresponding subgroups $\Lambda\subseteq\mathrm{PU}(2,1)$. Following work of Deligne–Mostow [DeligneMostow_BallQuotients], there has been a surge of interest in finding explicit complex geometric constructions of ball quotient varieties, see [Hunt_GeographyOfComplexSurfacesAnd3Folds] for an overview. Recently, a new example has been discovered by Deroin–Marché [DM_ToledoInvariantsTQFT, Section 6]. In this context, to show that a given variety is a ball quotient, the following criterion of Miyaoka proved very useful:
Theorem.
(Miyaoka [Miyaoka_ChernNumbersSurfacesGeneralType]) Let $X$ be a smooth complex projective surface such that $K_{X}$ is big and nef. Then $X$ is isomorphic to a quotient of the complex unit ball $\mathbb{B}^{2}$ if and only if
$3c_{2}(X)-c_{1}(X)^{2}=0.$
More generally, Yau [Yau_KEMetrics] proved that a smooth projective variety of dimension $n$ with ample canonical bundle $K_{X}$ is a ball quotient if and only if
(1) $\Big(2(n+1)c_{2}(X)-nc_{1}(X)^{2}\Big)\cdot K_{X}^{n-2}=0.$
Under the assumption that $K_{X}$ is merely big and nef, Greb–Kebekus–Peternell–Taji [GKPT_MY_Inequality_Uniformisation_of_Canonical_models] showed that if (1) holds, then the canonical model $X_{\mathrm{can}}$ is a, possibly singular, ball quotient variety. Note, however, that in this case $X$ itself may or may not be isomorphic to a ball quotient [Mul_Thesis]. The primary goal of this note is to extend [GKPT_MY_Inequality_Uniformisation_of_Canonical_models] to give a complete characterization of ball quotients among all minimal smooth projective varieties of general type, purely in terms of their Chern numbers:
Theorem A.
Let $X$ be a smooth complex projective variety of dimension $n$ such that $K_{X}$ is big and nef. Then $X$ is isomorphic to a quotient of the complex unit ball $\mathbb{B}^{n}$ if and only if
$\left((n+1)^{i}\cdot c_{i}(X)-\binom{n+1}{i}c_{1}(X)^{i}\right)\cdot K_{X}^{n-i}=0,\qquad\forall i=1,\ldots,n.$
The author is grateful towards Ben McKay for pointing out that Theorem A is intimately related to the existence of holomorphic normal projective connections. Indeed, if $X$ is a complex manifold which admits such a connection, then the series of equalities in Theorem A holds, see [KO80, Thm. 3.1]. Under this additional assumption, the conclusion that $X\cong\mathbb{B}^{n}$ was obtained previously by Jahnke–Radloff [JR02]. On the other hand, the assumptions in Theorem A are a priori much weaker than the existence of such a connection.
We will deduce Theorem A from the following more general result:
Theorem B.
Let $X$ be a smooth complex projective variety of dimension $n$ such that $K_{X}$ is big and nef. Assume, moreover, that
(2) $c_{i}(X)\cdot K_{X}^{n-i}=\frac{1}{(n+1)^{i}}\cdot\binom{n+1}{i}\cdot c_{1}(X)^{i}\cdot K_{X}^{n-i},\qquad\forall i=1,\ldots,k-1,$
for some integer $2\leq k\leq n$. Then
(3) $c_{k}(X)\cdot K_{X}^{n-k}\geq\frac{1}{(n+1)^{k}}\cdot\binom{n+1}{k}\cdot c_{1}(X)^{k}\cdot K_{X}^{n-k}.$
Moreover, the equality holds in (3) if and only if $X_{\mathrm{can}}\cong\mathbb{B}^{n}/\Lambda$ is isomorphic to a (possibly singular) ball quotient variety and the natural birational map
$f\colon X\rightarrow X_{\mathrm{can}}$
is an isomorphism over an open subset $U\subseteq X_{\mathrm{can}}$ with $\operatorname{codim}(X_{\mathrm{can}}\setminus U)\geq k$.
Note that (2) is always satisfied when $k=2$; in this case, Theorem B coincides precisely with [GKPT_MY_Inequality_Uniformisation_of_Canonical_models, Thm. 1.1, Thm. 1.2].
In recent years, inequalities between the first and second Chern class of projective varieties have been thoroughly investigated, see for example [GKT_OverviewUniformisation] or [Mul_Thesis, Introduction] for a comprehensive survey. However, little is known so far concerning the higher Chern classes. The boundedness of Chern classes of algebraic varieties of general type has been investigated in increasing generality by Kollár–Matsusaka [KollarMatsusaka_RRInequalities], Hunt [Hunt_GeographyOfComplexSurfacesAnd3Folds], Lu [LU_RRTypeInequalities], Du–Sun [DuSun_ChernClassInequalities], Ping–Zheng [PF_ChernCLassesPolarisedManifolds] and Lu–Xiao [LuXiao_InequalitiesOfRiemannRochType]; in particular, it is known [DuSun_ChernClassInequalities] that there exists a constant $C_{n}>0$ such that
$\left|\frac{c_{\lambda}(X)}{c_{1}(X)^{n}}\right|\leq C_{n}$
for any smooth projective variety $X$ of dimension $n$ such that $K_{X}$ (or $K_{X}$) is ample and any monomial Chern class of top degree $c_{\lambda}(X)$. However, the works above do not concern themselves with the question of finding the optimal value for $C_{n}$. Note that examples of Schreieder–Tasin [SchreiederTasin_ChernClassesUnbounded] show that the boundedness fails if one does not require positivity of the canonical bundle, even when fixing the underlying diffeomorphism type of $X$.
On the other hand, for threefolds $X$ fibred in minimal smooth projective surfaces, Lu–Tan–Zuo [LTZ_CanonicalClassInequalityFibredThreefolds] showed that $18c_{3}(X)>c_{1}(X)^{3}$, while for smooth projective threefolds with ample canonical bundle $\omega_{X}$, Sun [Sun_bogomolovtypeinequalitiesfrobenius] recently showed that
$ch_{3}(\Omega_{X}\otimes\omega_{X}^{\otimes 6})\leq\frac{19^{3}}{54}\cdot K_{X}^{3}.$
The author is unaware of any further research on the behaviour of the higher Chern classes of minimal varieties of general type. It seems interesting to further investigate this subject.
Throughout this text, we work over the field $\mathbb{C}$ of complex numbers. We employ the standard notions and conventions from [lazarsfeld_PositivityI].
Acknowledgments
I am grateful to thank Daniel Greb for many interesting discussions on Chern classes and for his support. Moreover, I would like to thank Marc Levine for suggesting me to learn about stringy invariants. The stimulative working environment and hospitality of Stefan Kebekus and the Universität Freiburg, where parts of this article were written, is much appreciated.
During the creation of this article, I was partially supported by the DFG research training group 2553 ‘Symmetries and classifying spaces: analytic, arithmetic and derived’.
2. Stringy Euler number
In this section, we recall the definition of the so-called stringy Euler number, which will be useful in proving Theorem B. The concept of stringy invariants first arose in physics [DHVW_StringsOnOrbifoldsI, DHVW_StringsOnOrbifoldsII, ZaslowStringyInvariants] and was later adapted by Batyrev [Batyrev_MotivicIntegrationStringyEulerNumbers] to the context of birational geometry: let $(X,D)$ be a log smooth pair and assume that $(X,D)$ is sub-klt. In other words, writing $D=\sum_{i=1}^{s}d_{i}\cdot D_{i}$, we assume that $d_{i}<1$. For any subset $I\subseteq\{1,\ldots,s\}$ denote
$D_{I}=\left\{\begin{matrix}\bigcap_{i\in I}D_{i},&I\neq\emptyset\\ X,&I=\emptyset\end{matrix}\right.\quad\mathrm{and}\quad D^{\circ}_{I}=D_{I}\setminus\bigcup_{i\in\{1,\ldots,s\}\setminus I}D_{I}\cap D_{i}.$
The stringy Euler number of $(X,D)$ is the rational number
$e_{\mathrm{Str}}(X,D):=\sum_{I\subseteq\{1,\ldots,s\}}\left(\prod_{i\in I}\frac{1}{1-d_{i}}\right)\ e(D_{I}^{\circ})$
Example 2.1.
Let $X$ be a smooth variety. Then $e_{\mathrm{Str}}(X):=e_{\mathrm{Str}}(X,0)=e(X)$, cf. [Fulton_ToricVarieties, p. 142].
The definition of stringy Euler number can be extended to arbitrary sub-klt pairs using the following fundamental Theorem of Batyrev:
Theorem 2.2.
(Batyrev [Batyrev_MotivicIntegrationStringyEulerNumbers, Thm. 1.5]) Let $\phi\colon(X_{1},D_{1})\rightarrow(X_{2},D_{2})$ be a proper, crepant birational morphism between log smooth sub-klt pairs. Then
$e_{\mathrm{Str}}(X_{1},D_{1})=e_{\mathrm{Str}}(X_{2},D_{2}).$
[…]
is called the stringy Euler number of $(X,D)$; by Theorem 2.2, it is well-defined and independent of the choice of $\phi$.
Strictly speaking, Batyrev only proves Theorem 2.2 in case the exceptional locus of $\phi$ is a divisor with simple normal crossings. The general case is readily reduced to this one by passing to a higher resolution. Alternatively, the statement of Theorem 2.2 follows from [FLNU_StringyChernClasses, Thm. 0.1].
3. Proof of main results
In this section, we prove Theorems A and B. The following Lemma, which is proved using the concept of stringy Euler numbers, is the key ingredient:
Lemma 3.1.
Let $S$ be a normal projective variety of dimension $n$ with isolated quotient singularities only. Assume that there exists a finite quasi-étale Galois cover $\pi\colon S^{\prime}\rightarrow S$ by a smooth variety $S^{\prime}$. Assume, moreover, that there exists a proper crepant birational morphism $f\colon Z\rightarrow S$ such that $Z$ is a smooth variety. Then
$c_{n}\left(Z\right)\geq c_{n}(S):=\frac{c_{n}\big(S^{\prime}\big)}{\deg\pi},$
with equality if and only if $S$ is smooth. In this case, $f\colon Z\rightarrow S$ is an isomorphism.
Proof.
Denote by $G$ the Galois group of $\pi\colon S^{\prime}\rightarrow S$ and set $d:=\deg\pi$. By the Poincaré–Hopf theorem [Milnor_TopologyDifferentialTopology, p. 35],
(3.1) $c_{n}\left(Z\right)=e\left(Z\right),\quad\mathrm{and}\quad c_{n}\left(S^{\prime}\right)=e\left(S^{\prime}\right).$
Now, by definition, $e\left(Z\right)=e_{\mathrm{Str}}(S)$. On the other hand,
$e_{\mathrm{Str}}(S)=e(S_{{\rm{reg}}})+\sum_{i=1,\ldots,s}e_{\mathrm{Str}}(U_{i}),$
where we denote $S^{{\rm{sing}}}=\{x_{1},\ldots,x_{s}\}$ and where $x_{i}\in U_{i}\subseteq S$ are some small (contractible) analytic neighbourhoods. Let $G_{i}\subseteq G$ be the isotropy group of $\pi\colon S^{\prime}\rightarrow S$ at $x_{i}$. Then, by [Batyrev_MotivicIntegrationStringyEulerNumbers, Thm. 1.9],
(3.2) $e_{\mathrm{Str}}(U_{i})=\#\Big\{\big\{hgh^{-1}\big|h\in G_{i}\big\}\Big|g\in G_{i}\Big\}=:c_{i}$
is the number of conjugacy classes of $G_{i}$ under the action on itself by conjugation. As $\pi$ is étale of degree $d$ over $S_{{\rm{reg}}}$,
(3.3) $e\big(S_{{\rm{reg}}}\big)=\frac{1}{d}\ e\Big(\pi^{-1}\big(S_{{\rm{reg}}}\big)\Big)=\frac{1}{d}\ \left(e(S^{\prime})-\sum_{i=1}^{s}\#\left(\pi^{-1}(x_{i})\right)\right)=\frac{e(S^{\prime})}{d}-\sum_{i=1}^{s}\frac{1}{\#G_{i}}.$
Combining (3.1), (3.2) and (3.3), we obtain the formula
$c_{n}\left(Z\right)=e\left(Z\right)=e_{\mathrm{Str}}(S)=e(S_{{\rm{reg}}})+\sum_{i=1}^{s}e_{\mathrm{Str}}(U_{i})=\frac{c_{n}\left(S^{\prime}\right)}{d}+\sum_{i=1}^{s}\left(c_{i}-\frac{1}{\#G_{i}}\right).$
Finally, observe that $c_{i}\geq 1$ and $G_{i}\geq 1$, with equality if and only if $G_{i}=1$ is trivial. Hence,
$c_{i}-\frac{1}{\#G_{i}}\geq 0,$
with equality if and only if $G_{i}=1$ is trivial. This, in turn, is equivalent to $\pi$ being étale over $x_{i}$, that is $x_{i}\in S$ being a smooth point. In summary,
$c_{n}\left(Z\right)\geq\frac{c_{n}\left(S^{\prime}\right)}{d}$
with equality if and only if $S$ is smooth. We conclude by noting that in the latter case $f$ is necessarily an isomorphism since a non-trivial birational morphism between smooth varieties can never be crepant, cf. [kollarMori_BirationalGeometry, Cor. 2.31.(2), Cor. 2.63]. ∎
Proof of Theorem B.
We proceed by induction on $k\geq 2$. The case $k=2$ is just [GKPT_HarmonicMetricsUniformisation]. Now, assume that $k>2$. According to [GKPT_HarmonicMetricsUniformisation], $B:=X_{\mathrm{can}}\cong\mathbb{B}^{n}/\Lambda$ is a, possibly singular, ball quotient variety. By the induction hypothesis, $f\colon X\rightarrow B$ is an isomorphism over an open subset with complement of codimension at least k1k-1, and we need to show that
(3.4) $c_{k}(X)\cdot K_{X}^{n-k}\geq\frac{1}{(n+1)^{k}}\binom{n+1}{k}\cdot c_{1}(X)^{k}\cdot K_{X}^{n-k},$
with equality if and only if $f\colon X\rightarrow B$ is an isomorphism over an open subset with complement of codimension at least $k$.
Indeed, choose a finite quasi-étale Galois cover $\pi\colon B^{\prime}\rightarrow B$ by a smooth ball quotient variety $B^{\prime}$, of degree $d$ say. Fix a sufficiently large integer $\ell\geq 1$ and general hypersurfaces $A_{1},\ldots,A_{n-k}\in|\ell K_{B}|$. Set
$S:=A_{1}\cap\ldots\cap A_{n-k},\quad Z:=f^{-1}(S),\quad\mathrm{and}\quad S^{\prime}:=\pi^{-1}(S).$
By construction, $Z$ is smooth and $f\colon Z\rightarrow S$ is a crepant birational morphism which is an isomorphism away from finitely many points $x_{1},\ldots,x_{s}\in S$. In particular, $S$ has at most finitely many isolated singularities. Moreover, $\pi\colon S^{\prime}\rightarrow S$ is a finite quasi-étale Galois cover, of degree $d$, by the smooth variety $S^{\prime}$.
Now, consider the short exact sequence
$0\rightarrow\mathcal{T}_{Z}\rightarrow\mathcal{T}_{X}|_{Z}\rightarrow\mathcal{N}_{Z/X}\rightarrow 0.$
Set $D_{i}:=f^{-1}(A_{i})$. Using that the total Chern class is multiplicative on short exact sequences we see that
$\begin{split}&\ell^{n-k}\cdot c_{k}(X)\cdot K_{X}^{n-k}=c_{k}(X)\cdot D_{1}\cdots D_{n-k}=c_{k}(X)|_{Z}\\ &\qquad=\sum_{i=0}^{k}c_{k-i}(Z)\cdot c_{i}\Big(\mathcal{N}_{Z/X}\Big)=\ell^{i}\cdot\binom{k}{i}\ (K_{X}|_{Z})^{i}.\end{split}$
Here, we also used that $\mathcal{N}_{Z/X}\cong\mathcal{O}(D_{1})\oplus\ldots\oplus\mathcal{O}(D_{m})$ to obtain the final equality. In any case, a similar argument shows that
$c_{k}(B^{\prime})\cdot K_{B^{\prime}}^{n-k}=\frac{1}{\ell^{n-k}}\sum_{i=0}^{k}\ell^{i}\cdot\binom{k}{i}\cdot c_{k-i}\big(S^{\prime}\big)\cdot(K_{B^{\prime}}|_{S^{\prime}})^{i}.$
As $f\colon Z\rightarrow S$ is an isomorphism away from finitely many points of $S$, we see that $d\cdot c_{k-i}(Z)\cdot(K_{X}|_{Z})^{i}=c_{k-i}(S^{\prime})\cdot(K_{B^{\prime}}|_{S^{\prime}})^{i}$ for all $i\geq 1$ and, consequently,
(3.5) $c_{k}(X)\cdot K_{X}^{n-k}-\frac{c_{k}(B^{\prime})\cdot K_{B^{\prime}}^{n-k}}{d}=\frac{1}{\ell^{n-k}}\left(c_{k}(Z)-\frac{c_{k}(S^{\prime})}{d}\right)\geq 0.$
Here, the inequality on the right is due to Lemma 3.1. On the other hand, since $B^{\prime}$ is a smooth ball quotient variety,
(3.6) $c_{k}\big(B^{\prime}\big)\cdot K_{B^{\prime}}^{n-k}=\frac{\binom{n+1}{k}}{(n+1)^{k}}\cdot c_{1}\big(B^{\prime}\big)^{k}\cdot K_{B^{\prime}}^{n-k}=\frac{d\cdot\binom{n+1}{k}}{(n+1)^{k}}\cdot c_{1}(X)^{k}\cdot K_{X}^{n-k}.$
Putting (3.5) and (3.6) together, we deduce that
$c_{k}(X)\cdot K_{X}^{n-k}\geq\frac{1}{(n+1)^{k}}\binom{n+1}{k}\cdot c_{1}(X)^{k}\cdot K_{X}^{n-k}.$
In case of equality, Lemma 3.1 shows that $f\colon Z\rightarrow S$ is an isomorphism. Equivalently, $f\colon X\rightarrow B$ is an isomorphism away from a subset of codimension $k$ in $B$, as required. ∎
Proof of Theorem A.
Follows immediately from Theorem B. ∎